OpenStudy (anonymous):

What is the magnitude of vector with ihat= -3, jhat= -9, and khat= 8?

OpenStudy (anonymous):

\[\hat i - 3 \; , \; \hat j -9 \;,\; \hat k\;\; 8\] ????

OpenStudy (anonymous):

...that doesn't seem to make sense to me

OpenStudy (anonymous):

it is \[\sqrt{(-3)^{2} + (-9)^{2}+(8)^{2}}\] which is 12.4 I think

OpenStudy (anonymous):

oh there was an equal sign...that made sense now

OpenStudy (anonymous):

The magnitude is equal to the square root of the (sum of the squares of the components) or: \[|A|=\sqrt{x^2 + y^2 + z^2} = \sqrt{ 9 + 81 + 64} = \sqrt{154} \approx 12.41\]

OpenStudy (anonymous):

To help make more sense think of a cube